OpenStudy (anonymous):

Can someone help me please? In a triangle EFG, EF=10, FG=8 and EG=14. Find the area of triangle EFG?

OpenStudy (anonymous):



OpenStudy (kc_kennylau):

Since it is not a right angled triangle, I suggest that you use the Heron's Formula

OpenStudy (kc_kennylau):

\[A=\sqrt{s(s-a)(s-b)(s-c)}\]where \(s=\dfrac{a+b+c}2\)

OpenStudy (anonymous):

do u mind substitueting the letters with the numbers in the example? so that i understand where they go etc.

OpenStudy (kc_kennylau):

a,b,c are the lengths of the sides

OpenStudy (kc_kennylau):

\[s=\frac{8+10+14}2=16\]

OpenStudy (kc_kennylau):

\[\mbox{Area}=\sqrt{16(16-8)(16-10)(16-14)}\\=\sqrt{16\times8\times6\times2}=\sqrt{16^2\times6}=16\sqrt6\]

OpenStudy (anonymous):

ok thats cool thxxx

OpenStudy (kc_kennylau):

So you understand the Heron's Formula now?